OpenStudy (anonymous):

A coffee shop mixes Brazilian coffee worth $5 per kilogram with Turkish coffee worth $8 per kilogram. The mixture is to sell for $7 per kilogram. How much of the Brazilian coffee should be used to make 300 kg of the mixture?

OpenStudy (anonymous):

B = kg of Brazillian......T = kg of Turkish B + T = 300 5B + 8T = 7(B + T)